OpenStudy (anonymous):

can someone help me? how do you ask for help? Candy selling at $6.00 per pound will be mixed with candy selling at $9.00 per pound. How many pounds of the more expensive candy are needed to produce a 15-pound mixture that sells for $7.00 per pound?

OpenStudy (anonymous):

@amistre64 i need help do I have to do something to get help? im not sure how to use openstudy

OpenStudy (amistre64):

you post and either wait or tag, when people have the time they will respond.

OpenStudy (amistre64):

what are your thoughts on this problem?

OpenStudy (anonymous):

i thought I had to write the equations as fractions to equal 15 but it doesnt work

OpenStudy (amistre64):

well lets talk our way thru it x pounds of candy1, added to y pounds of candy, have give us 15 lbs together is x + y = 15 a good start?

OpenStudy (amistre64):

now x pounds at $6 a pound costs 6x y pounds at $9 a pound costs 9y when combined, they should cost $7 6x + 9y = 7 is also a consideration to this problem

OpenStudy (anonymous):

I know its somewhere between 3 and 4 pounds of the $9 candy

OpenStudy (amistre64):

since x and y are the same, it is a solution to the system of both equation x+y = 15 6x + 9y = 7 now lets say x = 7-n, and y=8+n x = 7-n +y = 8+n ---------- x+y = 15 for all x and y so lets use that in the other one 6x + 9y = 7 6(7-n) + 9(8+n) = 7 now we can solve for n and thereby know x and y

OpenStudy (anonymous):

sorry I didnt see your post

OpenStudy (amistre64):

this is just basic a idea, a strategy (one of many), does it make sense tho?

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

3.56

OpenStudy (amistre64):

im not getting a realistic solution, at least not from the wolf.

OpenStudy (amistre64):

x+y = 15 6x + 9y = 7 either the problem has no solution, or ther eis something wrong with the setup here

OpenStudy (amistre64):

7(15) we have 15 pounds

OpenStudy (amistre64):

x+y = 15 6x + 9y = 7(15) thats better

OpenStudy (anonymous):

oh I got it but probably not the right way 10 pounds of the $6 candy and 5 pounds of the $9 candy 60+45= 105 divided by 15 so it would be 7 not the right way to do it

OpenStudy (anonymous):

your way is the right way

OpenStudy (amistre64):

there is no 'right' or wrong way .. there are multiple methods that can reach the same conclusion, as long as we have the correct information to play with :) 10 of x, and 5 of y is fine

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

hat math i have dyscalculia so any math is really hard for me to compute and takes me longer

OpenStudy (anonymous):

I hate :P

OpenStudy (amistre64):

i just old and forgetful lol

OpenStudy (anonymous):

lol thanxs